# A CMC existence result for expanding cosmological spacetimes

### v2

@@ 4.1 Extending Theorem 3 to the case $\Lambda>0$
Ric= $\text{Ric}\,=\,8\pi\left(T-\frac{1}{n-1}\text{tr}_{g}(T)g\right)+\frac{2}{n-1}\Lambda g.$ (4.10)
By imposing a standard condition on the energy-momentum tensor [21, Section 4.3], one is led to the following version of the strong energy condition in the case of a positive cosmological constant,
@@
$\frac{d\mathcal{H}}{du}\leq\lambda-\mathcal{H}^{2},$
where $\mathcal{H}=\frac{H}{n}$. The theorem again follows from standard comparison results (see e.g., [23, Section 1.6]), [12, Section 2]). ∎
@@ 4.2 Remarks on extending the results in [11] to the case $\Lambda>0$
We show how [11, Prop. 4.6] generalizes to the case when (4.11) is assumed. However, in this new setting, the constant $C$ appearing in the statement of [11, Prop. 4.6] will now also depend on $\lambda$. From Proposition 3.3(ii) and eq. (1) in [11], we have the following equation
@@
where $\mathscr{H}$ is the forcing term and $\Delta_{s}$ is the spatial Laplacian induced from the Riemannian metric on $V_{s}$. (Note there is a typo in the proof of [11, Prop. 4.6].)
@@
Assuming (4.11) and the monotonicity assumption on $\mathscr{H}$ (as assumed in [11], in particular as in our case in which $\mathscr{H}=c=$ const.), we have
@@
$-f^{2}H^{2}\,=\,-f^{2}(H-\mathscr{H}+\mathscr{H})^{2}\,\leq\,-f^{4}+2|f^{3}\mathscr{H}|\,.$
Using the Peter-Paul inequality twice, we estimate the cubic term:
@@
[…]
@@
The rest of the proof proceeds exactly as in the proof of [11, Prop. 4.6].
@@ 5 Further comments on Conjecture 2
In this last section, we provide some further comments on Conjecture 2 and its relationship to the future causal boundary. In section 5.1 we discuss some issues concerning the structure of the future causal boundary of cosmological spacetimes. In section 5.2 we establish a connection between the future causal boundary being spacelike and the existence of a maximal Cauchy surface. The relationship to the Bartnik splitting conjecture is also discussed.
@@ 5.1 Some remarks on the causal boundary of cosmological spacetimes
Let $(M,g)$ be a spacetime which contains a compact Cauchy surface and obeys the strong energy condition, ${\rm Ric}(X,X)\geq 0$ for all timelike vectors $X$. If $(M,g)$ is timelike geodesically complete, then $(M,g)$ splits isometrically into the product (×$({\mathbb{R}}\times V,-dt^{2}\oplus h)$, where $(V,h)$ is a compact Riemannian manifold.
@@
We make a few remarks about the proof of Theorem 1. The proof involves the notion of the future causal boundary of a spacetime. Let $(M,g)$ be a globally hyperbolic spacetime. Very briefly, the future causal boundary $\mathscr{C}^{+}$ consists of the set of all terminal indecomposable past sets, or TIPs for short. These are the sets of the form $I^{-}(\gamma)$, where $\gamma$ is a future inextendible timelike curve in $M$, with appropriate identifications, when two such curves have the same timelike past. See e.g., [21, Section 6.8], for a more detailed discussion.
@@
where the $a_{i}$’s are positive, and the $x^{i}$’s should be viewed as periodic coordinates. It is convenient to assume each $x^{i}$ lies in an interval, $-b\leq x^{i}\leq b$, with the end points identified. Basic causal theoretic arguments show that the tori, $V_{t}=\{t\}\times T^{3}$ are Cauchy surfaces, and hence that $(M,g)$ is globally hyperbolic.
@@
It is not difficult to show by direct computation that the resulting model satisfies the strong energy condition, in fact it satisfies ${\rm Ric}(X,X)\geq 0$ for all vectors $X$. (Alternatively, one could apply [9, Prop. 3.2].) As a hint to this, observe that,
@@
[…]
From the $x^{3}$-translational symmetry, a similar conclusion to the above holds for each $t$-line centered at $x^{1}=x^{2}=0,x^{3}=\xi$, $-b\leq\xi\leq b$ ($b>4$), and, moreover, each such $t$-line defines a distinct element of $\mathscr{C}^{+}$. It can be further shown that $\mathscr{C}^{+}$ consists only of these elements. In fact, much of this follows from a general result of Harris [20, Prop. 3.5], the relevant part of which we now paraphase.
@@
where $(a,b)$ is an interval (possibly $a=\infty$ or $b=\infty$), and where for $i=1,...,m$, $(V_{i},h_{i})$ is a Riemannian manifold.
The following is immediate from [20, Prop. 3.5].
@@
This result applied to our example (5.13), (5.14) implies that $\mathscr{C}^{+}$ is homeomorphic to $S^{1}$, as is somewhat suggestive by the discussion above. Furthermore, $\mathscr{C}^{+}$ is spacelike. In general, the future causal boundary being spacelike may be taken to mean that no TIP is properly contained in another TIP.
Thus, the simple example presented here, which can be generalized in various directions, shows that, under the assumptions of Conjecture 2, the future causal boundary need not consist of a single element. One, however, may be inclined to formulate a new question: Under the assumptions of Conjecture 2, will the future causal boundary always be spacelike? An example showing otherwise could not, of course, be within this class of multiply warped product spacetimes. Moreover, an affirmative answer would settle the Bartnik splitting conjecture. For, as shown in [18], in this case the spacetime admits a timelike line, and one can apply the Lorentzian splitting theorem under the assumptions of the Bartnik splitting conjecture. In the next section we provide an alternative proof of this fact. We show that if a spacetime satisfies the hypotheses of Conjecture 2 but is also past timelike complete and has a spacelike future causal boundary, then the spacetime contains a maximal Cauchy surface and consequently splits [5].
@@ 5.2 Existence of a maximal Cauchy surface from a spacelike causal boundary
Suppose $(M,g)$ satisfies the assumptions of Conjecture 2 but is also past timelike complete. If $M$ satisfies the ray-to-ray condition, then $M$ contains a maximal Cauchy surface.
A spacelike future causal boundary implies the ray-to-ray condition, see [18, Prop. 5.9]. Thus we have the following corollary.
@@
$S^{-}_{\infty}\,=\,\partial\left(\bigcup_{k=0}^{\infty}I^{-}(S^{-}_{k})\right),$
@@
We briefly allude to the proof of these properties. $S^{-}_{\infty}$ is an achronal boundary by construction and hence it’s edgeless. Moreover, $S^{-}_{\infty}$ does not intersect $I^{+}(S_{2})$ and so $S^{-}_{\infty}\subset J^{-}(S_{2})$; consequently, $S^{-}_{\infty}$ is a past Cauchy surface by [17, Prop. 3.17] and acausal by [17, Thm. 3.18]. Lastly, (4) follows from a similar argument as in Proposition 4, see [17, Thm. 4.2].
Similarly, let $S^{+}_{\infty}$ be the future ray horosphere associated to $\eta|_{[t_{1},\infty)}$. It also satisfies analogous properties of (1) - (4) above. Consider the compact set $K:=J^{+}(S_{1})\cap J^{-}(S_{2})$. Using (3) above and the analogous property for $S^{+}_{\infty}$, it’s not hard to see that
@@
Consequently, $J^{-}(S^{-}_{\infty})\cap J^{+}(S^{+}_{\infty})$ is compact. The existence of a maximal Cauchy surface now follows from [13, Thm. 3.7]. The proof of this theorem is somewhat involved, so we give a more direct proof.
Since $J^{+}(S^{+}_{\infty})\cap J^{-}(S^{-}_{\infty})$ is compact, there exist points $x_{1}\in S^{+}_{\infty}$ and $x_{2}\in S^{-}_{\infty}$ such that $\tau:=d(x_{1},x_{2})$ realizes the Lorentzian distance between $S^{+}_{\infty}$ and $S^{-}_{\infty}$. Now set $\Sigma=\{x\in M\mid d(x,S^{+}_{\infty})=\tau\}$; using the fact that $S^{+}_{\infty}$ is a future Cauchy surface, arguments as in the proof of Proposition 4 show that $\Sigma$ is a closed acausal $C^{0}$ hypersurface, and moreover it lies locally to the future of $S^{-}_{\infty}$ near $x_{2}$. By [13, Lem. 3.6] $\Sigma$ has mean curvature $H\leq 0$ in the support sense. (The proof involves, modulo a certain technicality, pushing the support hypersurfaces of $S^{+}_{\infty}$ along the normal geodesics and making use of the Raychaudhuri equation.) We are now in a position to apply a rough version of the Lorentzian geometric maximum principle, [3, Thm. 3.6], from which we conclude that $\Sigma$ and $S^{-}_{\infty}$ intersect in a smooth spacelike maximal hypersurface containing $x_{2}$. A continuity argument shows that this intersection globalizes. Thus $\Sigma=S^{-}_{\infty}$ is a smooth maximal hypersurface which is closed and acausal. Since it’s contained in the compact set between $S_{1}$ and $S_{2}$, it is also compact and consequently a Cauchy surface. ∎
@@ Acknowledgments
We thank Piotr Chruściel, Stacey Harris, Gerhard Huisken, and Alex Mramor for helpful comments and discussions. Gregory Galloway was supported by the Simons Foundation, Award No. 850541. Eric Ling was supported by Carlsberg Foundation CF21-0680 and Danmarks Grundforskningsfond CPH-GEOTOP-DNRF151. Part of the research on this paper was performed while the authors were participants of the program “New Frontiers in Curvature: Flows, General Relativity, Minimal Submanifolds, and Symmetry” at the Simons Laufer Mathematical Sciences Institute (SLMath). We thank SLMath for its support.

### v3

@@ 4.1 Extending Theorem 3 to the case $\Lambda>0$
$\text{Ric}\,=\,8\pi\left(T-\frac{1}{n-1}\text{tr}_{g}(T)g\right)+\frac{2}{n-1}\Lambda g.$ (4.10)
By imposing a standard condition on the energy-momentum tensor [22, Section 4.3], one is led to the following version of the strong energy condition in the case of a positive cosmological constant,
@@
$\frac{d\mathcal{H}}{dr}\leq\lambda-\mathcal{H}^{2},$
where $\mathcal{H}=\frac{H}{n}$. The theorem again follows from standard comparison results (see e.g., [24, Section 1.6]), [13, Section 2]). ∎
@@ 4.2 Remarks on extending the results in [12] to the case $\Lambda>0$
We show how [12, Prop. 4.6] generalizes to the case when (4.11) is assumed. However, in this new setting, the constant $C$ appearing in the statement of [12, Prop. 4.6] will now also depend on $\lambda$. From Proposition 3.3(ii) and eq. (1) in [12], we have the following equation
@@
where $\mathscr{H}$ is a smooth function on $M$ and $\Delta_{s}$ is the spatial Laplacian induced from the Riemannian metric on $V_{s}$. (Note there is a typo in the proof of [12, Prop. 4.6]. A parenthesis is misplaced in the first line of the proof.)
@@
Assuming (4.11) and the monotonicity assumption as in [12], $\langle\nabla\mathscr{H},\nu\rangle\geq 0$ (which is trivially satisfied in our situation $\mathscr{H}=c=$ const.), we have
@@
$f^{2}H^{2}\,=\,-f^{2}(H-\mathscr{H}+\mathscr{H})^{2}\,\leq\,-f^{4}+2|f^{3}\mathscr{H}|\,.$
Using the Peter-Paul inequality ($ab\leq\tfrac{a^{2}}{2\varepsilon}+\tfrac{\varepsilon b^{2}}{2}$) twice, we estimate the cubic term:
@@
$2n\lambda f^{2}\,\leq\,2n\lambda\left(\frac{f^{4}}{2\varepsilon_{3}}+\frac{\varepsilon_{3}}{2}\right).$
@@
This inequality is equivalent to the last inequality in the proof of [12, Prop. 4.6], where the constant now also depends on $\Lambda$. The rest of the proof is as in [12].
@@ 5 Further comments on Conjecture 2
In this last section, we once again assume the strong energy condition. We provide some further comments on Conjecture 2 and its relationship to the future causal boundary. In section 5.1 we discuss some issues concerning the structure of the future causal boundary of cosmological spacetimes. In section 5.2 we establish a connection between the future causal boundary being spacelike and the existence of a maximal Cauchy surface. The relationship to the Bartnik splitting conjecture is also discussed.
@@ 5.1 Some remarks on the causal boundary of cosmological spacetimes
Let $(M,g)$ be a spacetime which contains a compact Cauchy surface and obeys the strong energy condition, ${\rm Ric}(X,X)\geq 0$ for all timelike vectors $X$. If $(M,g)$ is timelike geodesically complete, then $(M,g)$ splits isometrically into the product $({\mathbb{R}}\times V,-dt^{2}\oplus h)$, where $(V,h)$ is a compact Riemannian manifold.
@@
We make a few remarks about the proof of Theorem 1. The proof involves the notion of the future causal boundary of a spacetime. Let $(M,g)$ be a globally hyperbolic spacetime. Very briefly, the future causal boundary $\mathscr{C}^{+}$ consists of the set of all terminal indecomposable past sets, or TIPs for short. These are the sets of the form $I^{-}(\gamma)$, where $\gamma$ is a future inextendible timelike curve in $M$, with appropriate identifications, when two such curves have the same timelike past. See e.g., [22, Section 6.8], for a more detailed discussion.
@@
where the $a_{i}$’s are positive, and the $x^{i}$’s should be viewed as periodic coordinates. (Such spacetimes are sometimes used as spatially homogeneous models in cosmology [22].) It is convenient to assume each $x^{i}$ lies in an interval, $b\leq x^{i}\leq b$, with the end points identified. Basic causal theoretic arguments show that the tori, $V_{t}=\{t\}\times T^{3}$ are Cauchy surfaces, and hence that $(M,g)$ is globally hyperbolic.
@@
It is not difficult to show by direct computation that the resulting model satisfies the strong energy condition, in fact it satisfies ${\rm Ric}(X,X)\geq 0$ for all vectors $X$. (Alternatively, one could apply [10, Prop. 3.2].) As a hint to this, observe that,
@@
where $e_{0}=\frac{\partial}{\partial t}$. Moreover, it is not difficult to show that $M$ is future timelike geodesically complete; this essentially relies on the fact that $\int_{t_{0}}^{\infty}a_{i}dt=\infty$, $i=1,2,3$. (For a general completeness result applicable to multiply warped product spacetimes, see [29, Section 4].)
Now we consider the structure of the future causal boundary $\mathscr{C}^{+}$. Let $t\to\gamma(t)$ be the $t$-line centered at $x^{i}=0$, $i=1,2,3$. Let $s\to\eta(s)$ be a past directed null geodesic starting at $(t_{0},0,0,0)$ on $\gamma$, with $t_{0}>1$. With the choice $a_{3}=t^{\frac{5}{4}}$, it is not difficult to show that the intersection of every such null geodesic $\eta$ with the toroidal time slice $V_{1}=\{t=1\}$ is confined to the region $4<x^{3}<4$. This boils down to the fact that $\int_{1}^{t_{0}}t^{-\frac{5}{4}}dt<4$. Since $\partial I^{-}(\gamma(t_{0}))$ is ruled by such geodesics, it follows that $I^{-}(\gamma(t_{0}))$ does not meet the region $|x^{3}|\geq 4$. Since $I^{-}(\gamma)=\cup_{t_{0}\in(1,\infty)}I^{-}(\gamma(t_{0}))$, we conclude that $I^{-}(\gamma)\cap V_{1}$ does not meet the region $|x^{3}|\geq 4$. In particular $I^{-}(\gamma)\neq M$. This immediately implies that $\mathscr{C}^{+}$ consists of more than one element. However, one can say somewhat more.
From the $x^{3}$-translational symmetry, a similar conclusion to the above holds for each $t$-line centered at $x^{1}=x^{2}=0,x^{3}=\xi$, $b\leq\xi\leq b$ ($b>4$), and, moreover, each such $t$-line defines a distinct element of $\mathscr{C}^{+}$. It can be further shown that $\mathscr{C}^{+}$ consists only of these elements. In fact, much of this follows from a general result of Harris [21, Prop. 3.5], the relevant part of which we now paraphase.
@@
where $(a,b)$ is an interval (possibly $a=$-$\infty$ or $b=\infty$), and where for i=1,,mi=1,...,m, $(V_{i},h_{i})$ is a Riemannian manifold.
The following is immediate from [21, Prop. 3.5].
@@
This result applied to our example (5.13), (5.14) implies that $\mathscr{C}^{+}$ is homeomorphic to $S^{1}$, as is somewhat suggested by the discussion above. Furthermore, $\mathscr{C}^{+}$ is spacelike. In general, the future causal boundary being spacelike may be taken to mean that no TIP is properly contained in another TIP.
Thus, the simple example presented here, which can be generalized in various directions, shows that, under the assumptions of Conjecture 2, the future causal boundary need not consist of a single element. One, however, may be inclined to formulate a new question: Under the assumptions of Conjecture 2, will the future causal boundary always be spacelike? An example showing otherwise could not, of course, be within this class of multiply warped product spacetimes. Moreover, an affirmative answer would settle the Bartnik splitting conjecture. For, as shown in [19], in this case the spacetime admits a timelike line (i.e., a globally maximizing inextendible timelike geodesic), and one can apply the Lorentzian splitting theorem under the assumptions of the Bartnik splitting conjecture. In the next section we provide an alternative proof of this fact. We show that if a spacetime satisfies the hypotheses of Conjecture 2 but is also past timelike complete and has a spacelike future causal boundary, then the spacetime contains a maximal Cauchy surface and consequently splits [5].
@@ 5.2 Existence of a maximal Cauchy surface from a spacelike causal boundary
Suppose $(M,g)$ satisfies the assumptions of Conjecture 2 but is also past timelike complete. If $M$ satisfies the ray-to-ray condition, then $M$ contains a maximal (i.e. mean curvature zero) Cauchy surface.
A spacelike future causal boundary implies the ray-to-ray condition, see [19, Prop. 5.9]. Thus we have the following corollary.
@@
$S^{-}_{\infty}\,=\,\partial\left(\bigcup_{k=1}^{\infty}I^{-}(S^{-}_{k})\right),$
@@
We briefly allude to the proof of these properties. $S^{-}_{\infty}$ is an achronal boundary by construction and hence it is edgeless. Moreover, $S^{-}_{\infty}$ does not intersect $I^{+}(S_{2})$ and so $S^{-}_{\infty}\subset J^{-}(S_{2})$; consequently, $S^{-}_{\infty}$ is a past Cauchy surface by [18, Prop. 3.17] and acausal by [18, Thm. 3.18]. Lastly, (4) follows from a similar argument as in Proposition 4, see [18, Thm. 4.2].
Similarly, let $S^{+}_{\infty}$ be the future ray horosphere associated to $\eta|_{[t_{1},\infty)}$. It also satisfies analogous properties of (1) - (4) above. Consider the compact set $K:=J^{+}(S_{1})\cap J^{-}(S_{2})$. Using (3) above and the analogous property for $S^{+}_{\infty}$, it is not hard to see that
@@
Consequently, $J^{-}(S^{-}_{\infty})\cap J^{+}(S^{+}_{\infty})$ is compact. The existence of a maximal Cauchy surface now follows from [14, Thm. 3.7]. The proof of this theorem is somewhat involved, so we give a more direct proof.
Since $J^{+}(S^{+}_{\infty})\cap J^{-}(S^{-}_{\infty})$ is compact, there exist points $x_{1}\in S^{+}_{\infty}$ and $x_{2}\in S^{-}_{\infty}$ such that $\tau:=d(x_{1},x_{2})$ realizes the Lorentzian distance between $S^{+}_{\infty}$ and $S^{-}_{\infty}$. Now set $\Sigma=\{x\in M\mid d(x,S^{+}_{\infty})=\tau\}$; using the fact that $S^{+}_{\infty}$ is a future Cauchy surface, arguments as in the proof of Proposition 4 show that $\Sigma$ is a closed acausal $C^{0}$ hypersurface, and moreover it lies locally to the future of $S^{-}_{\infty}$ near $x_{2}$. By [14, Lem. 3.6] $\Sigma$ has mean curvature $H\leq 0$ in the support sense. (The proof involves, modulo a certain technicality, pushing the support hypersurfaces of $S^{+}_{\infty}$ along the normal geodesics and making use of the Raychaudhuri equation.)
Hence to summarize, there is a spacetime neighborhood $U$ of $x_{2}$ such that $\Sigma$ lies to the future of $S^{-}_{\infty}$ within $U$ and $\Sigma$ has mean curvature $H\leq 0$ in the support sense, whereas $S^{-}_{\infty}$ has mean curvature $H\geq 0$ in the support sense. It then follows from a rough version of the Lorentzian geometric maximum principle, [3, Thm. 3.6], that there is an open set $A\subset\Sigma\cap S^{-}_{\infty}$ which is a smooth spacelike hypersurface with mean curvature H=0H=0.666Proposition 3.5 in [3] guarantees that, in our situation, a certain technical condition in the statement of [3, Thm. 3.6] is satisfied. A continuity argument then shows that this intersection globalizes. Thus $\Sigma=S^{-}_{\infty}$ is a smooth maximal hypersurface which is closed and acausal. Since it is contained in the compact set between $S_{1}$ and $S_{2}$, it is also compact and consequently a Cauchy surface. ∎
@@ Acknowledgments
We thank Piotr Chruściel, Stacey Harris, Gerhard Huisken, and Alex Mramor for helpful comments and discussions. Gregory Galloway was supported by the Simons Foundation, Award No. 850541. Eric Ling was supported by Carlsberg Foundation CF21-0680 and Danmarks Grundforskningsfond CPH-GEOTOP-DNRF151. Part of the research on this paper was supported by the National Science Foundation under Grant No. DMS-1928930 while the authors were in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Fall 2024 semester.